16. A 100 gm sample of steam starts at 100°C. How much heat in joules is lost condensing the steam into hot water? (a) 2.0x106 joules (b) 2.0x104 joules (c) 2.26-10, joules (d) 2.26-107 joules (e) 2.26지0, joules 17. A 100 gm sample of hot water starts at 100°C, & then loses 21,851 joules of heat. What is the final temper of the hot water? (a) 60.7°C (b) 52.2°C (c) 47.8°C (d) 95.2°C (e) 59.9°C
Answer #1

Q16. Enthalpy of evaporation or heat lost from steam at 100 C to water at 100 C = 2257 kj/kg = 2.26 x 106 / 1000 g =2.26 x 105 /100 g.

so answer is option c

Q17. Let the final temperatue = T

so heat lost = mass x specific heat x (100 - T)

= 100 x 4.186 x (100 - T) = 21851

so, T = 100-52.2= 47.8 C

so answer is option c

all the best in the course work